the difference between the squares of two consecutive integers is 133. what are the integers
Answer by ewatrrr(24785)

 
Hi,
the difference between the squares of two consecutive integers is 133.
Let x and (x+1) represent the two consecutive integers
Question states***
(x +1)^2 - x^2 = 133
x^2+2x+1-x^2 = 133
2x+1 = 133
2x = 132
x = 66 | Integers are 66,67
CHECKING our Answer***
4489 - 4356 = 133